OpenStudy (flvskidd):

A cyclist rides her bike at a speed of 24 kilometers per hour. What is this speed in kilometers per minute? How many kilometers will the cyclist travel in 15 minutes? Do not round your answers.

OpenStudy (evoker):

24km / hr equals \[\frac{ 24km }{ 1 hr}\]

OpenStudy (evoker):

in general you want to use a conversion fraction that changes the bottom from hours to minutes.

OpenStudy (evoker):

There are 60min in an hour.

jhonyy9 (jhonyy9):

you can using the rule of 3 simple than you know that rides 24 km per hour what mean 24 km per 60 minutes so this mean 24 km ------- 60 min. x km -------- 1 min. --------------------- x = 1*24/60 = 24/60 = ? hope helped

jhonyy9 (jhonyy9):

do you understand it ?

OpenStudy (mathmale):

Start with \[\frac{ 24km }{ 1 hr}\] Because 1 hour = 60 minutes, multiply \[\frac{ 24km }{ 1 hr}\] by \[\frac{ 1hr }{ 60 \min}\]

OpenStudy (mathmale):

Show your work, please. this will give you the speed in km/min.